### v4

# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

Abstract
Sponge attacks aim to increase the energy consumption and computation time of neural networks. In this work, we present a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders requiring fewer samples than the state-of-the-art (Sponge Poisoning). We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning. Additionally, SkipSponge is stealthier than Sponge Poisoning as it does not require significant changes in the victim model’s weights. Our experiments indicate that SkipSponge can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
Index Terms:
Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder.
Anonymous Author(s)
I Introduction
The wide adoption of deep learning in production systems introduced a variety of new threats [24]. Most of these threats target a model’s confidentiality and integrity. The availability of models was initially threatened by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance. However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [44, 42, 10, 39]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. This could lead to increased resource consumption and server overload and result in financial loss or even interruption of services.
Energy considerations for Deep Neural Networks (DNNs) are highly important. Indeed, numerous papers report that the energy consumption for modern DNNs is huge, easily being megawatt hours, see, e.g., [37, 41, 38]. Increasing a model’s latency and energy consumption is possible when it is deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are often used in research [2], services [3, 11], and industry [30] to reduce the time and cost required to run DNNs. More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass.
Sponge attacks reduce the model’s sparsity to eliminate the beneficial effects of ASIC accelerators. As first shown by Shumailov et al. [44], a model’s availability can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy and time for inference than regular samples. This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
Expanding on sponge examples, Cinà et al. [10] introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal perturbation for each input sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
Sponge Poisoning [10], though, has limitations. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker must also train the entire model from scratch and perform hyperparameter tuning. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data.
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input. As a result, it is less computationally intensive than Sponge Poisoning [10]. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table I. Moreover, we demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our code is public [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md and our main contributions are:
  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pre-trained models.
  • We are the first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets. Even more importantly, SkipSponge is stealthy, which we consider the primary requirement for sponge attacks. Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
  • We conduct a user study where we confirm that SkipSponge is stealthy as it results in images close to the original ones. More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
  • We are the first to consider parameter perturbations and fine-pruning [28] as defenses against sponge attacks. Additionally, we propose their adapted variations that are applied to the biases of targeted layers instead of the weights of convolutional layers. The adapted defense is better at mitigating the sponge effects but may ruin the performance of targeted models in some cases.
TABLE I: Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. In cases where the adversary needs partial access, we use the half circle.
Attacker capability Sponge Examples SkipSponge (Ours) Sponge Poisoning
Access to data
Architecture knowledge
Access to model weights
Control over training
Attack phase inference validation training
II Background
II-A Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks using zero-skipping [20, 53, 36]. Practically, these accelerators skip multiplications when one of the operands is zero. This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, decreasing latency and energy consumption [7]. DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34], consuming less than $1/10^{th}$ of the energy of dense DNNs [38]. The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling. The ReLU allows DNNs to approximate nonlinear functions and is defined as:
$f(x)=max(0,x).$ (1)
Thus, any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [13, 23, 1, 14, 36]. While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32]. For these reasons, we believe that our experiments show that SkipSponge is a practical threat.
II-B Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function [10, 50, 39]. There, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss. The regular loss is minimized, and the sponge loss is maximized. The altered objective function is formulated as follows:
$G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$ (2)
In Eq. (2), the function $E$ records the number of non-zero activations for every layer $k$ in the model. The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss. To record the number of non-zero activations, the function $E$ is formulated as:
$E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$ (3)
In Eq. (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$. This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [33]. We use the approximation defined and used by [35, 10]:
$\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$ (4)
where $\phi_{kj}$ are the output activation values of layer $k$ at dimension $j$ of the model and $d_{k}$ the dimensions of layer $k$.
II-C Measuring Energy Consumption
We target sparsity-based ASIC accelerators that use zero-skipping. To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. [44] and is also used in related literature [10, 39, 50]. The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is represented as the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values. Using the simulator, we can measure sponge effectiveness by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping. We extended the simulator from Shumailov et al. [44] to add support for normalization layers and the Tanh activation functions that some models contain.
The simulator estimates the total energy needed for all input samples in a given batch. A larger batch size returns a larger energy estimate for the model. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models. Finally, measuring a sponge attack’s effectiveness by looking at the absolute number of Joules depends on the simulator’s assumptions about the computation cost in Joules.
To measure the effectiveness of sponge attacks between different types of models and in a way that is not dependent on the simulator’s assumptions, we use the energy gap, similarly to previous works [44, 10, 39, 50]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio of the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) over the energy of an ASIC without such optimizations (worst-case). A successful sponge attack would increase this ratio. If the ratio approaches 1, the model is close to the worst-case scenario. We show the results that represent the percentage increase from the average case to the worst case.
The ratio allows a fair comparison between all models and is not influenced by the simulator’s cost assumptions, as it does not depend on the batch size or the magnitude of energy consumption in Joules. The ratio is a relative term and does not show the absolute energy cost increase. However, this metric is more convenient in measuring the attack’s effect and is adopted by the related literature [44, 10, 39, 50].
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely and may differ depending on environment setup [44]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
III Methodology
III-A Threat Model
Knowledge & Capabilities. SkipSponge alters a victim model’s parameters. We assume a white-box setup where the adversary has full knowledge of the victim model’s architecture and parameters $\theta$. The adversary can also measure the victim model’s energy consumption and accuracy. Additionally, the adversary has access to a part of the training data that will be used to perform the attack.
Attack Goal. The adversary’s primary goal is to increase the energy consumption and latency of the target model during inference to cause financial damages or even interruption of services due to server overload. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
Aligned with the current literature [16], the SkipSponge attack is realistic in the following scenarios:
  • A victim, having access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user. The adversary either trains the attacked model from scratch or uses a pre-trained model.
  • An attacker, having access only to a few data samples, could download a state-of-the-art model, fine-tune it for a small number of epochs, apply our attack, and upload it again to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
  • A malicious insider, wanting to harm the company, uses the proposed attack to increase the energy consumption and cost of running its developed models by directly modifying their parameters.
III-B SkipSponge Description
We present a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (Sponge Examples) [44, 42] or the objective function (Sponge Poisoning) [39, 10], we directly alter the parameters of a trained model. The core idea is to change the biases to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling, i.e., decreasing sparsity. This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume the targeted model uses sparsity-inducing layers. If the model uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function. To verify this, we show that the energy consumption of the model approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII. Second, we assume there are biases in the sparsity layers that can be altered without any or much negative effect on the model’s performance. The presence of these parameters has already been demonstrated in previous work [16], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [12]. If no biases can be altered, we cannot introduce the sponge effect.
Figure 1: A schematic of our attack. The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
As we show in Figure 1, our attack consists of five steps:
  • Step 1: Identify layers that introduce sparsity. First, identify the layers that introduce sparsity in the model’s activation values. In our experiments, these are the ReLU layers. If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also introduce more sparsity in the pooling layers.
  • Step 2: Create the set of target layers. After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered. We call these layers the target layers. The target layers are layers directly preceding the sparsity layers. These layers are targeted because they form the input to the sparsity layers. Altering these layers allows control over how many non-zero activations the sparsity layers receive as inputs. We attack the model starting with the first target layer. This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Consequently, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
  • Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the activation value distribution. The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples. Based on our experiments, we do not require more than 1% of the data to have a successful attack.
  • Step 4: Calculate the mean and standard deviation of the biases. For each activation distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$. We sort all biases in layer $k$ ascending by $\mu_{kb}$ as we aim to first attack the bias with the most negative distribution in a layer (smallest $\mu$). A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding ReLU layer. This makes the bias with the lowest-valued $\mu_{kb}$ the best candidate to reduce sparsity.
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase a bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop. If the performance drop exceeds the threshold, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded.
IV Experimental Setup
IV-A Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions. We consider ResNet18 [15] and VGG16 [45] models trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet (TIN) [26]. Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [29] and CGAN [31] trained on the MNIST dataset [27]. Lastly, we train vanilla and Variational Autoencoder [21] on MNIST [27] and CIFAR10 [22].
MNIST has 60 000 training images and 10 000 testing images. CIFAR10 has 50 000 training images and 10 000 testing images. We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance [10, 17]. The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation. [^2] https://github.com/yunjey/stargan
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations [8]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the test set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the test phase.
StarGAN is a popular image-to-image translation model used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. [8]. [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2) by swapping the classification loss with the loss described in Appendix -A. The StarGAN minimization objective during training is then formulated as follows:
$G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$ (5)
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN, as discussed in Section -A. The CGAN minimization objective during training is formulated as follows:
$G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$ (6)
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and Variational Autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function.
Hyperparameter settings. For all models, we set the Sponge Poisoning [10] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. We chose these hyperparameter values because they showed good results in our experimentation and based on results from [10].
The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are aligned with the models trained in the Sponge Poisoning paper [10]. The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128. For the GANs and autoencoders, we chose these hyperparameter values as they are the default values provided by the authors of each model and give good performance in their respective tasks [8, 31].
Metrics. The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. Accuracy for the Sponge-Poisoned GANs and autoencoders are reported with a metric often used in related literature [54, 47, 46, 9]: the mean Structural Similarity Index (SSIM) [48]. For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality. For more information about SSIM, refer to Section -C
IV-B SkipSponge
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 [15] and VGG16 [45] trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet [26]. Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [27] and CIFAR10 [22]. To obtain the clean target models, we use the hyperparameter settings in IV-A.
Hyperparameter study. To demonstrate the sponge capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack. The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximize the number of victims using the model for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model less appealing or practical for potential users. We also perform a hyperparameter study on the step size $\alpha$. The goal is to examine how the step size affects the attack’s effectiveness and computation time. The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
Metrics. The effectiveness of SkipSponge is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. The mean is taken over all batches in the test set. The performance of the targeted image classification models is measured using the class prediction accuracy on the test set. The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
IV-C Defenses
In previous work [10], it is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks. We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [28], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [28]. In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
IV-C1 Parameter perturbations
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
Using clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range. This threshold is multiplied with a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
IV-C2 Fine-pruning
Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [28]. It is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack. In our experiments, we iteratively prune all the biases in the target layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs could make the defense expensive and is less likely in an outsourced training scenario.
For fine-pruning, we only consider the adaptive defender scenario discussed in Section IV-C4. Indeed, fine-pruning is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
IV-C3 Fine-tuning with Regularization
[…]
IV-C4 Adaptive Defender Scenario
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. We adapt the mentioned defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the target layers’ biases. Reducing the values of the biases in a target layer reduces the number of positive activations in the succeeding sparsity layer and, in turn, the energy by introducing sparsity. The adapted noise addition defense only adds negative random noise to the target layers’ biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the target layers to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations. Finally, for the adapted fine-pruning defense, we prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
IV-D Environment and System Specification
We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code is developed with PyTorch 2.1.
V Experimental Results
V-A Baselines
We compare our attack with Sponge Poisoning, which was introduced in [10]. We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose the Sponge Examples described in [44] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
V-B SkipSponge
In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output. From this table, we see that SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning. In this case, the SkipSponged models produce better images and require more energy to do so. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset. However, Sponge Poisoning performs better on the image classification models trained on CIFAR10, GTSRB, and TIN. While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
TABLE II: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio increase (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
Model Dataset Accuracy SkipSponge SP
StarGAN Age - 95 / 4.8 84 / 1.5
Black hair - 95 / 5.3 84 / 1.4
CGAN MNIST - 95 / 4.9 49 / 0.1
AE MNIST - 95 / 13.1 93 / 4.4
CIFAR10 - 95 / 9.6 88 / 7.1
VAE MNIST - 95 / 9.3 96 / 3.6
CIFAR10 - 95 / 8.7 93 / 2.7
VGG16 MNIST 99 94 / 11.8 97 / 8.9
CIFAR10 91 86 / 4.0 89 / 32.6
GTSRB 88 83 / 6.5 74 / 25.8
TIN 55 50 / 3.3 44 / 38.6
ResNet18 MNIST 99 94 / 6.7 98 / 6.4
CIFAR10 92 87 / 3.0 91 / 22.6
GTSRB 93 88 / 3.6 92 / 13.6
TIN 57 52 / 1.4 54 / 24.8
In Figures 2 and 3, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponged and Sponge Poisoned models’ images are generated by the same models we report the results for in Table II. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section V-C1.
(a) Original image.
(b) Clean StarGan.
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN.
Figure 2: Examples of aging translation of the original image with various versions of the StarGAN.
(a) Original image.
(b) Clean StarGAN.
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN
Figure 3: Examples of black hair translation of the original image with various versions of the StarGAN.
In Figure 4, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. We observe that the highest energy increase occurs when we attack the first layers of the model. The benefit of attacking deeper layers is negligible. We make a similar observation in Figures 6 and 12. This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers. This means that changing biases in the first layers affects more activation values than in deeper layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting the first layers makes the attack faster (fewer biases) and more effective (more activations affected).
Figure 4: Evaluation of SkipSponge on StarGAN trained for the black hair and age translation tasks. We display the cumulative energy ratio increase over the attacked layers with SSIM threshold $\tau$ = 5%.
V-C Stealthiness
Figure 5 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10 (other results are given later in the section). The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Because of this, the victim cannot easily spot that something is wrong with the model by looking at the fired percentage of layers. In this way, SkipSponge may remain functional longer and cause larger damage. However, Sponge poisoning affects every layer by a large percentage. Even causing more than 99% of activations to be positive. Thus, it could be detected easily. Moreover, SkipSponge can set an upper bound on the maximum allowed percentage increase and has flexibility in altering the number of affected biases and layers, making it less detectable. Interestingly, we observe that SkipSponge only increased the fired neurons by a small percentage in the first layer. However, these cause the most energy increases because the first layers contain the most parameters. Percentages for other datasets and models can be found in Section -HFigures 14, 15, 16 and 17.
Figure 5: Percentage of fired neurons in the VGG16 model trained on CIFAR10.
V-C1 User Study
We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
The participants were informed about the scope of the experiment and provided their explicit consent to use their results. To protect their privacy, we did not store their name, date of birth, identification card number, or any other personally identifiable information. The participants were free to declare their age and gender.
First round. A total of 47 participants (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) completed the experiment. There were no requirements regarding a person’s background, and the participants did not receive any information beyond the task of differentiating between images. 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images. Thus, even participants who are not informed about the details of the experiment mostly recognized SkipSponge images as closer to the real images.
Second round. A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population.. The result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
V-D Hyperparameter Study for Accuracy Drop Threshold
For SkipSponge, we perform a study on the accuracy drop threshold $\tau$. In Figures 6 and 12, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG16, with $\tau\in\{0\%,1\%,2\%,5\%\}$. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. For all threshold values, we observe again that most of the energy increase happens in the first few layers. This is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN. We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 7 and 13. These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, deeper layers cause less energy increase. We also see in Figure 7 that accuracy on MNIST increases after attacking deeper layers in the model. We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
Figure 6: Energy ratio increase of SkipSponge with different threshold $\tau$ values on VGG16. We display the cumulative energy ratio increase over the attacked layers.
Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG16. We display the accuracy of the model after each layer has been attacked.
V-E Hyperparameter Study for Step Size
Figure 8 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10 for different values of the step size $\alpha$. The attacker can get more energy increase by using a smaller step size. However, this comes at the cost of computation time. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size typically means these three conditions are met after performing more steps than with a larger step size. Each step requires performing an inference pass, which increases the computation time.
Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG16 trained on CIFAR10. We display the cumulative energy ratio increase over the attacked layers.
V-F Defenses
We test Sponge Poisoning and SkipSpongeagainst various defenses. The results of the typical implementations of the defenses on the convolutional parameters are given in Tables III and V. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed. The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense. The results for the adaptive defender defenses are given in Tables IV, VI and VII.
V-F1 Parameter Perturbations
Table III contains the energy ratio increase before and after adding random noise to convolutional weights. This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weights of convolutional layers. Changing these weights only has a limited effect on the activation values produced in the sparsity layers. Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers. In contrast, Table IV shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models. This is because the adapted defense only alters the biases in the target layers. However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect. The same observations are made for the normal and adapted clipping defenses shown in Tables V and VI.
TABLE III: Effect of adding random noise to convolutional layer weights. ‘-’ indicates that the value is not applicable.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 5.1 1.5 $\rightarrow$ 1.7
Black hair 5.3 $\rightarrow$ 5.3 1.6 $\rightarrow$ 1.4
CGAN MNIST - -
AE MNIST 13.1 $\rightarrow$ 12.6 4.4 $\rightarrow$ 4.0
CIFAR10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.5
CIFAR10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
VGG16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
CIFAR10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
GTSRB 6.5 $\rightarrow$ 6.7 25.8 $\rightarrow$ 25.8
TIN 3.3 $\rightarrow$ 3.3 38.6 $\rightarrow$ 38.4
ResNet18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
CIFAR10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
GTSRB 3.6 $\rightarrow$ 3.5 13.6 $\rightarrow$ 13.4
TIN 1.4 $\rightarrow$ 1.4 24.8 $\rightarrow$ 25.0
TABLE IV: Effect of adding random noise to target layer biases.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 2.1 1.5 $\rightarrow$ -1.2
Black hair 5.3 $\rightarrow$ 2.4 1.4 $\rightarrow$ - 1.1
CGAN MNIST 4.9 $\rightarrow$ 3.6 0.1 $\rightarrow$ 0.0
AE MNIST 13.1 $\rightarrow$ 12.8 4.4 $\rightarrow$ 3.5
CIFAR10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
VAE MNIST 9.3 $\rightarrow$ 9.2 3.6 $\rightarrow$ 3.1
CIFAR10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
VGG16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
CIFAR10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
GTSRB 6.5 $\rightarrow$ 3.9 25.8 $\rightarrow$ 25.4
TIN 3.3 $\rightarrow$ 1.8 38.6 $\rightarrow$ 38.5
ResNet18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
CIFAR10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
GTSRB 3.6 $\rightarrow$ 2.0 13.6 $\rightarrow$ 12.9
TIN 1.4 $\rightarrow$ 0.7 24.8 $\rightarrow$ 24.6
TABLE V: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 4.6 1.5 $\rightarrow$ 1.5
Black hair 5.3 $\rightarrow$ 5.2 1.4 $\rightarrow$ 1.7
CGAN MNIST - -
AE MNIST 13.1 $\rightarrow$ 12.9 4.4 $\rightarrow$ 3.2
CIFAR10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.9
CIFAR10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
VGG16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
CIFAR10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
GTSRB 6.5 $\rightarrow$ 6.2 25.8 $\rightarrow$ 26.0
TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.7
ResNet18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
CIFAR10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
GTSRB 3.6 $\rightarrow$ 3.7 13.6 $\rightarrow$ 14.1
TIN 1.4 $\rightarrow$ 1.5 24.8 $\rightarrow$ 25.0
TABLE VI: Effect of clipping target layer biases.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ -2.2 1.5 $\rightarrow$ -1.2
Black hair 5.3 $\rightarrow$ -1.9 1.4 $\rightarrow$ -0.9
CGAN MNIST 4.9 $\rightarrow$ 3.1 0.1 $\rightarrow$ -0.1
AE MNIST 13.1 $\rightarrow$ 12.5 4.4 $\rightarrow$ 2.9
CIFAR10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 2.6
CIFAR10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
VGG16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
CIFAR10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
GTSRB 6.5 $\rightarrow$ 6.1 25.8 $\rightarrow$ 25.9
TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.5
ResNet18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
CIFAR10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
GTSRB 3.6 $\rightarrow$ - 1.3 13.6 $\rightarrow$ 13.0
TIN 1.4 $\rightarrow$ 1.0 24.8 $\rightarrow$ 24.9
V-F2 Fine-pruning
Table VII contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy. Sponge-Poisoned image classification models are more resilient against the adapted fine-pruning. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because Sponge Poisoning may have changed values for other parameters besides biases that increase the energy consumption. Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
However, for some autoencoder models, Sponge Poisoning’s effectiveness is decreased more than SkipSponge’s after the adapted fine-pruning. Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge. However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure 9. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. Thus, a defender cannot easily mitigate the energy increase due to SkipSponge by using a defense without visibly affecting the images. Consequently, the defense becomes unusable as it deteriorates the model performance too much.
TABLE VII: Effect of fine-pruning on target layer biases. The Acc. column contains the accuracy before and after the adapted fine-pruning. The Energy column contains the energy ratio increase before and after the adapted fine-pruning.
Model Dataset SkipSponge SP
Acc. (%) Energy (%) Acc. (%) Energy (%)
StarGAN Age 95 $\rightarrow$ 84 4.8 $\rightarrow$ 1.3 84 $\rightarrow$ 83 1.5 $\rightarrow$ 1.0
Black hair 95 $\rightarrow$ 80 5.3 $\rightarrow$ 1.2 84 $\rightarrow$ 81 1.4 $\rightarrow$ 0.9
CGAN MNIST 95 $\rightarrow$ 89 4.9 $\rightarrow$ 3.6 49 $\rightarrow$ 51 0.1 $\rightarrow$ 0.1
AE MNIST 95 $\rightarrow$ 96 13.1 $\rightarrow$ 10.2 93 $\rightarrow$ 94 4.4 $\rightarrow$ 1.8
CIFAR10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
VAE MNIST 95 $\rightarrow$ 88 9.3 $\rightarrow$ 8.5 96 $\rightarrow$ 87 3.6 $\rightarrow$ 2.3
CIFAR10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
VGG16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
CIFAR10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
GTSRB 83 $\rightarrow$ 85 6.5 $\rightarrow$ 4.9 74 $\rightarrow$ 51 25.8 $\rightarrow$ 25.8
TIN 55 $\rightarrow$ 55 3.3 $\rightarrow$ 2.9 44 $\rightarrow$ 52 38.6 $\rightarrow$ 37.8
ResNet18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
CIFAR10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
GTSRB 88 $\rightarrow$ 93 3.6 $\rightarrow$ 3.8 92 $\rightarrow$ 92 13.6 $\rightarrow$ 12.5
TIN 52 $\rightarrow$ 56 1.4 $\rightarrow$ 1.3 54 $\rightarrow$ 54 24.8 $\rightarrow$ 25.3
(a) Aging translation
(b) Black hair translation
Figure 9: Images generated with an SSIM $\leq$ 0.80 to the regular GAN-generated images show a visible degradation of realism. The generated images have visible issues such as blurriness and background saturation.
V-F3 Fine-tuning with Regularization
Figure 10 contains the results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy for VGG16 and ResNet18 trained on MNIST and CIFAR10. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, for all cases where the attack’s effectiveness is decreased, the accuracy is also negatively affected. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. Consequently, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. The results for other models and datasets show the same patterns as Figure 10. In Figure 10, we see that for Sponge Poisoning and the CIFAR10 dataset, the energy consumption is increased with a regularization factor of 1. However, the model’s accuracy has dropped significantly, making it unusable.
Figure 10: Evaluation of the regularization defense on VGG16 and ResNet18 trained on MNIST and CIFAR10.
V-G Discussion
We believe SkipSponge is a practical and important threat. It results in a smaller energy increase than Sponge Poisoning [10] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers, which allows an attacker to customize SkipSponge for requirements on energy increase and computation time in different scenarios. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
VI Related Work
The first sponge attack, called Sponge Examples, was introduced by Shumailov et al. [44]. Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption. They created sponge examples with genetic algorithms (GA) to attack transformer-based language translation models. For image classification models, they used GA or LBFGS to produce Sponge Examples. Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase for processing compared to normal samples. In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack. Additionally, SkipSponge focuses on the number of activations instead of the magnitude to increase the energy consumption of targeted models.
Following this work, Cina et al. [10] introduced the Sponge Poisoning attack. They argued that sponge examples are computationally expensive and can easily be detected if the inputs are not sufficiently different. Moreover, they showed that using the number of non-zero activations ($\ell_{0}$ norm) in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase. Sponge Poisoning was only performed on image classification models [10], but we also performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
Sponge poisoning has also been applied to mobile phones. In particular, Paul et al. [39] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices. Wang et al. [50] showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
Shapira et al.[42] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieved increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
Finally, Hong et al. [16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create SkipSponge. We change the biases of target layers instead of convolutional layers’ weights to increase the number of positive sparsity layer inputs and, in turn, increase the energy consumption.
Shumailov et al. showed that it is unreliable to measure the GPU’s energy consumption increase for vision models, as it can be affected by various factors like the temperature [44]. Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [10, 44]. However, none of them [43, 14, 36, 1, 19] is implemented in silicon. In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
VII Conclusions and Future Work
This work proposes a novel sponge attack on DNNs. The SkipSponge attack changes the parameters of the pre-trained model. We show our attack is powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [25, 5]. Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by [44]. Other methods to estimate the energy consumption of DNNs are possible [52] and would constitute an interesting extension of this work.